OpenStudy (sunshinehaggard):

Imagine that you are standing on the ground 15 feet away from a flagpole. If the top of the flag is 25 feet high, which equation represents the angle θ your feet make with the top of the flag?

OpenStudy (mathstudent55):

A drawing may be a big help.

OpenStudy (sunshinehaggard):

this is all it gives me and the choices and i dont understand

OpenStudy (mathstudent55):

Draw a figure. Try it and show what you did.

OpenStudy (08surya):

tan(theta)=25/15=5/3